GlueX TOF Meeting, March 18, 2011

From GlueXWiki
Jump to: navigation, search

Location, Time

Friday, March 18, 2011
2:00 pm EST
JLab: CEBAF Center, Room F326


  1. Announcements
  2. Minutes from the last meeting
  3. Prototype Status
  4. Light guides
    1. Design
  5. Detector Rates and PMT performance
  6. Contract Status



  1. ESNet: 8542553
  2. EVO: EVO site


  1. dial:
  2. enter access code followed by the # sign: 3421244#


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2011-1Q on the JLab CUE. This directory is accessible from the web at .



  • FSU: Sasha Ostrovidov
  • JLab: Eugene Chudakov, Mark Ito (chair), Elton Smith, Sascha Somov, Simon Taylor, Tim Whitlatch, Beni Zihlmann

Review of minutes from the last meeting

We went over the minutes from the March 4th meeting.

  • Sasha O. received the new F1 TDC that Beni sent to FSU. He sees the same set of multiple peaks with the new one that he observed with the old one. In addition both TDC's are seeing hits in channels that are not connected to anything. Sascha S. had a couple of recommendations:
    1. Send email to Ed Jastrzembski with symptoms.
    2. Look at the pulse width, verify that it is at the expected 50 ns.
  • FSU is still waiting for the 25-ps-least-count TDC from CAEN. See Paul Eugenio's email.
    • Sasha S.[?] pointed out that the CAEN TDC needs to be looked at from DAQ point of view, i. e., clock distribution, trigger, etc.

Light guides

Beni pointed out the wiki page he is putting together to collect information on the light guide design. It summarizes a lot of the discussions we have had over the past few meetings and includes relevant drawings of conceptual designs.

Elton thought that we probably should try to simulate light collection for the various ideas. He has a program that will do that; he agreed to help Beni get started using it.

Detector rates and PMT performance

Beni showed another wiki page he is that addresses the problem if high rates on the innermost, long TOF counters. Sascha S. reminded us of the note that he wrote that estimates these rates for the TOF as well as for other detectors.

Sasha O. reported that at one point, due to a light leak, he observed rates of around 5 MHz in one of the counters. Signal characteristics at that high a rate is a question.

Elton thought that to reduce the load on these PMT's we may have to have amplifiers on them so that they can be run at lower gain. Eugene thought that perhaps the amplifiers could be put into the PMT base. Fernando will be consulted about possible designs.

Contract Status

Mark showed a draft Advanced Procurement Plan (APP). This gives a schedule for completion of a construction contract for the TOF between FSU and JLab. He hopes to get it signed next week.

Action Items

  • Consult with JLab electronics experts on TDC problems. -> Sasha O.
  • Get started with light guide simulation. -> Elton, Beni
  • Link to Sascha S.'s rate note from Beni's rate wiki page. -> Beni
  • Ask Fernando about amplifiers for inner counters. -> ?
  • Comprehensive report on geometry options at a future meeting. -> Paul
  • Get APP signed. -> Mark